simplify the expression (x^8y^2/x^-2y^-8)^1/2 i feel like this may not make sense (it could just be me)i haven't learned fractional exponents yet
Answer by richwmiller(17219)

x^(1/2) means the sqrt(x)
y^(-8) mean 1 over y^8 =1/(y^8)
x^-2=1/(x^2)
After dealing with all the negative exponents we get
sqrt((x^8 y^2)/(y^8/x^2))
the x^2 gets put in the numerator because it is in the denominator of the denominator
sqrt((x^10/(y^6))
taking the sqrt gives us x^5/y^3
So the whole thing comes down to x^5/y^3